Question 28
ProbabilityThree men, P, Q and R aim at a target the probabilities that P, Q and R hit the target are ½, ⅓ and ¾ respectively. Find the probability that exactly two of them hit the target.
- A.1
- B.½
- C.5/12
- D.⅓
Show the answer and explanation
Answer: C (5/12)
Explanation
Add the three cases with one miss: (½)(⅓)(¼) + (½)(⅔)(¾) + (½)(⅓)(¾) = 1/24 + 6/24 + 3/24 = 10/24 = 5/12.